OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). m + n = 1000 0.05m + 0.06n = 57

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

m + n = 1000 (1) 0.05m + 0.06n = 57 (2) frm (1) we get m = 1000 -n substitute this in (2) 0.05(1000 - n) + 0.06n = 57 50 - 0.05n + 0.06n = 57 0.01n = 57 - 50 n = 7/0.01 n = 700

OpenStudy (anonymous):

now substitute n = 700 in (1) m + 700 = 1000 m = 1000 - 700 m = 300 so solution set is (300, 700)

OpenStudy (anonymous):

if u r satisfied with the answer, pls click GOOD ANSWER button for me....☺

OpenStudy (anonymous):

but what is the answer for x any y

OpenStudy (anonymous):

normally we use x and y as variables and while writing the solution we first write the value of x then y in the form (x,y) in your case, the variables are m and n instead of x and y so (x,y) is same as (m,n) = (300, 700)